The Once-ler produces Thneeds (which everyone needs) in a factory that also produces pollution. The more pollution the Once-ler produces, the lower his production costs, with the marginal benefit of x to the Once-ler equal to MBx = 50–2x, where x is the amount of pollution produced. This pollution harms the local population of Brown Bar-Ba-Loots according to the following marginal cost function: MCx = 20+3x.

21. How much pollution will the Once-ler produce if left to his own devices?

a) 17

b) 25

c) 50

d) 6

e) None of the above

22. In a landslide, The Lorax is elected president. What is the optimal quantity regulation that The Lorax would set?

a) 50

b) 6

c) 25

d) 17

e) None of the above
Answer #1

21)

Without any regulation, firm will produce pollution such that MB=0

50-2x=0

x=25

Correct option is

b) 25

22)

Optimal level of pollution can be achieved if MB=MC i.e.

50-2x=20+3x

5x=30

x=6

Lorax will set a regulation that limits the pollution to 6 units

Correct option is

b) 6